2 months ago Michael deposited R9000 into savings account at a simple interest rate of 11.5% per year. He also deposited R5000 six months ago into another savings account earning 8% per year compounded half yearly. The total amount of Michael's savings in 3 years time is

1
Expert's answer
2021-11-24T18:19:14-0500

Simple interest


"A=P(1+rt)"


Time ("t)" is "3 years +\frac{2}{12} months"


"A = 9000(1+(0.115*3.167))"


"A =" "R12,277.85"


Compound interest


"A = P(1+\frac{r}{n})^{nt}"


Time ("t") is "3 years + \frac{6}{12} months"


"A = 5000(1+\frac{0.08}{2})^{2*3.5}"


"A = R 6,579.66"


Total amount


"TA =" "R12,277.85" "+" "R6,579.66"


"TA = R18,857.51"